There are three possible values of the random variable Y representing the number of tails. These are 0, 1, and 2. Assign probability values P(Z) to each value of the random variable

1
Expert's answer
2022-03-07T08:07:06-0500

1. Let's denote "H" - heads, "T" - tails.

There are "2^2=4" possible outcomes


"S=\{HH, HT, TH, TT\}"

The possible values of the random variable "Y" are "0, 1, 2."

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Possible \ Outcomes & Y \\ \hline
 HH & 0 \\
 \hdashline
 HT & 1 \\
 \hdashline
 TH & 1 \\
 \hdashline
 TT & 2 \\
\end{array}"


We will assume that the probability of getting heads and tails is the same:


"p = q =1\/2"


Construct the probability distribution of the random variable


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 y & 0 & 1 & 2 \\ \hline
 p(y) & 1/4 & 1/2 & 1/4
\end{array}"

or


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 y & 0 & 1 & 2 \\ \hline
 p(y) & 0.25 & 0.5 & 0.25 
\end{array}"